OpenStudy (jackfrost33):

The value of x in the proportion 1/2:2/3 = 3/4:x is 1 4/9 1 7/9 1/4

OpenStudy (anonymous):

1/2:2/3 is the same as (1/2) / (2/3). And as you probably know by now, dividing by a fraction is the same as multiplying by its reciprocal So (1/2) / (2/3) = (1/2) * (3/2) = 3/4 For what value of x will 3/4 = 3/4:x? Well, x = 1 :)

OpenStudy (jackfrost33):

thx